OpenStudy (anonymous):

The diameter of a soccer ball is 8.65 inches. What is the volume of the ball, rounded to the nearest hundredth? A. 313.42 in3 B. 323.18 in3 C. 338.88 in3 D. 234.06 in3

OpenStudy (anonymous):

\[Volume of sphere = 4/3 \pi r^3\]

OpenStudy (anonymous):

1/2 Diameter = Radius (r)

OpenStudy (anonymous):

Can you solve it with this info now?